OpenStudy (zzr0ck3r):

show that cosh(t) = (e^t+e^(-t))/2

OpenStudy (anonymous):

Given what definition of \(\cosh(t)\)?

OpenStudy (zzr0ck3r):

geomtric using hyperbola

OpenStudy (anonymous):

that is?

OpenStudy (anonymous):

What are we given about \(\cosh(t)\) that we can start out with?

OpenStudy (zzr0ck3r):

im not quite sure, she drew a hyperbola and then said the area of a slice would be t/2 where t is the archlength then she said something 1/2cosh(t)sinh(t) - integral from 1 to cosh(t) sqrt(x^2-1) ....

OpenStudy (zzr0ck3r):

im not quite sure what was going on...

OpenStudy (zzr0ck3r):

one sec im finding my notes

OpenStudy (anonymous):

Me neither. I just know \(\large \cosh(t) = \frac{e^x+e^{-x}}{2}\) by definition

OpenStudy (zzr0ck3r):

these are my notes but they suck because I was lost, this was the first time I ever heard of cosh...

OpenStudy (anonymous):

Okay... so if we wanted to find the area, I'd recommend integrating with respect to y, because it is y simple.

OpenStudy (zzr0ck3r):

lol im lost

OpenStudy (anonymous):

Hmmm, never mind I messed up...

OpenStudy (zzr0ck3r):

I have a table that has the integral in it, I just dont understand what Im solving for, a budy of mine told be to sub x for cosh(h) in that integral I posted but I dont get why...

OpenStudy (anonymous):

okay can you give me the integral then?

OpenStudy (anonymous):

What does \(t\) represent? is it the angle?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (zzr0ck3r):

yeah one sec integral of sqrt(u^2-a^2) = u/2sqrt(u^2-a^2) - (a^2/2)ln(a+sqrt(u^2-a^2)) + C

OpenStudy (zzr0ck3r):

i think at one point I have to use series expansion of e^x i think...

OpenStudy (zzr0ck3r):

thats (u/2)*sqrt(u^2-a^2) - (a^2/2)*ln(a+sqrt(u^2-a^2)) + C

OpenStudy (anonymous):

Hmmm, okay so actually if I go back to my integration via y idea: 

OpenStudy (anonymous):

To make this a little less complicated: